A Bibliography of Papers in Lecture Notes in Computer Science (2009)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.03

Title word cross-reference

α [2]. -sensitivity [2]. 6th [6]. 9th [7, 8]. Algorithm [3]. algorithmics [9]. Algorithms [6]. Barcelona [6]. based [2]. Baton [7, 8]. Characterization [5]. China [9]. Computational [7, 8]. conference [7, 8].

Crawler [1]. Crawling [1]. Degree [5]. Dependence [5]. Early [1]. Efficiently [1]. equilibria [2]. FAW [9]. February [6]. Fractional [1]. Frontiers [9]. games [2]. Graph [3, 6]. Heat [3]. Hefei [9]. Homogeneous [4]. ICCS [7, 8]. II [8]. Important [1]. In-Degree [5]. Index [4]. international [7, 8, 6, 9]. June [9].

Kernel [3].

Linear [4]. Local [3].

May [7, 8]. models [6].

Nash [2]. network [2].

One [4].

PageRank [1, 2, 3, 4, 5]. PageRank-based [2]. Pages [1]. Part [7, 8]. Partitioning [3]. Prioritizing [1]. Proceedings [7, 8, 9, 6].

reputation [2]. Rouge [7, 8].

science [7, 8]. sensitivity [2]. Singular [4]. Spain [6]. Splitting [4]. Systems [4].

Tail [5]. Third [9].

URLs [1]. USA [7, 8]. Using [3].

WAW [6]. web [6]. web-graph [6]. workshop [6, 9].

References

Alam:2009:FPC

[1] Md. Hijbul Alam, Jongwoo Ha, and Sangkeun Lee. Fractional PageRank crawler: Prioritizing URLs efficiently for crawling important pages early. In Anonymous, editor, Proceedings of the 14th International Conference on Database Systems for Advanced Applications, volume 5463 of Lecture Notes in Computer Science, pages 590–594. Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 3-642-00886-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

Chen:2009:SNE

[2] Wei Chen, Shang-Hua Teng, Yajun Wang, and Yuan Zhou. On the α-sensitivity of Nash equilibria in PageRank-based network reputation games. In Deng et al. [9], pages 63–73. ISBN 3-642-02269-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/
content/p1040n9jt618;   http://www.
zentralblatt-math.org/zmath/en/search/
?an=1166.68003.

Chung:2009:LGP

[3] Fan Chung. A local graph partitioning algorithm using heat kernel PageRank. In Avrachenkov et al. [6], pages 62–75. ISBN 3-540-95994-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

Jager:2009:PSH

[4] Douglas V. Jager and Jeremy T. Bradley. PageRank: Splitting homogeneous singular linear systems of index one. In Anonymous, editor, Proceedings of the 2nd International Conference on Theory of Information Retrieval: Advances in Information Retrieval Theory, volume 5766 of Lecture Notes in Computer Science, pages 17– 28. Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 3-642-04416-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

Litvak:2009:CTD

[5] Nelly Litvak, Werner Scheinhardt, Yana Volkovich, and Bert Zwart. Characterization of tail dependence for in-degree and PageRank. In Avrachenkov et al. [6], pages 90–103. ISBN 3-540-95994-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

Avrachenkov:2009:AMW

[6] Konstantin E. Avrachenkov, Debora Donato, and Nelly Litvak, editors. Algorithms and models for the web-graph: 6th international workshop, WAW 2009 Barcelona, Spain, February 12–13, 2009 proceedings, volume 5427 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 3-540-95995-5, 3-540-95994-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 W39 2009.

Allen:2009:CSIa

[7] Gabrielle Allen, Jarosław Nabrzyski, Edward Seidel, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational science – ICCS 2009. 9th international conference Baton Rouge, LA, USA, May 25–27, 2009. Proceedings, Part I, volume 5544 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009.

Allen:2009:CSIb

[8] Gabrielle Allen, Jarosław Nabrzyski, Edward Seidel, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational science – ICCS 2009. 9th international conference Baton Rouge, LA, USA, May 25–27, 2009. Proceedings, Part II, volume 5545 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009.

Deng:2009:FAT

[9] Xiaotie Deng, John E. Hopcroft, and Jinyun Xue, editors. Frontiers in algorithmics. Third international workshop, FAW 2009, Hefei, China, June 20–23, 2009. Proceedings, volume 5598 of Lecture Notes in Computer Science.

Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 3-642-02270-7, 3-642-02269-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43 F39

                                        2009. URL http://www.springerlink.
                                        com/content/p1040n9jt618;   http://
                                        www.zentralblatt-math.org/zmath/en/
                                        search/?an=1166.68003.